40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 20 21:43:25 2010, yuta, Update, CDS, checking whitening filters for MCs and messed up MC2ULSEN.pngMC2LRSEN.png
    Reply  Thu Oct 21 21:35:16 2010, yuta, Update, CDS, fixed whitening filter switches LLLR.pngContec23BO.png
    Reply  Fri Oct 22 00:59:01 2010, yuta, Update, CDS, new MEDM screens, new filter banks(whitening and dewhitening related) HappySideMEDM.png
       Reply  Fri Oct 22 23:29:48 2010, yuta, Update, CDS, fixed output filter switching 
          Reply  Mon Nov 1 23:31:05 2010, yuta, Update, CDS, checked input hardware filter in single frequency 
Message ID: 3757     Entry time: Thu Oct 21 21:35:16 2010     In reply to: 3748
Author: yuta 
Type: Update 
Category: CDS 
Subject: fixed whitening filter switches 

(Joe, Yuta)

Summary:
 We found some mistakes in whitening filter switches yesterday(see elog #3748).
 One was the mis-connection between LL/LR and the second was the bit invert.
 So, we fixed it and checked that they are now switching correct by probing MAX333A.

Why we made mistake:
1. LL/LR mis-connection
  Simply, Simulink model was wrong.
  It occurred because UL/UR/LR/LL order is currently arbitrary and confusing.
  For now, we just swap the connection(Attached #1).
  In the future, we should fix all orders corresponding to the actual wiring order(UL > LL > UR > LR).

2. Bit invert
  It occurred from counter-intuitive output of Contec 32 BO(See Attachment #2(taken from this wiki page)).
  If "1," current goes from A to B, so MAX333A input will be 0V, which means analog WF is on.
  If "0," no current, so MAX333A input will be +15V, which means analog WF is off(bypassed).

Next work:
 - There are no digital WF in SDSEN inputs. So, we have to put them.
 - Dewhitening filters are totally messed up now. Switches are wrong, some optics have digital DW but some doesn't, some have 28Hz elliptic LPF.......
     We have to grand unify them.

Attachment 1: LLLR.png  8 kB  Uploaded Thu Oct 21 20:10:58 2010  | Hide | Hide all
LLLR.png
Attachment 2: Contec23BO.png  48 kB  Uploaded Thu Oct 21 22:41:41 2010  | Hide | Hide all
Contec23BO.png
ELOG V3.1.3-